# Large Language Models for Blockchain Security: A Systematic Literature Review

### v2

@@ 3 Taxonomy of LLM4BS task
In this section, we introduce a thematic taxonomy devised to systematically categorize the body of literature about tasks associated with large language models for blockchain security (LLM4BS), emphasizing the function of the LLM within these contexts. Fig. 1 depicts the five applications of LLM4BS task.
@@ 3.1 LLM as Code auditor on Smart Contracts
The application of LLM in the domain of smart contract code auditing and vulnerability detection can be succinctly encapsulated as follows: Advanced tools Wang et al. (2024); Sun et al. (2024); David et al. (2023); Karanjai et al. (2023); Sönmez and Knottenbelt (2024); ORTU et al. (????); Sun et al. (2023); Yu et al. (2023); Sun et al. (2024) powered by Large Language Models, such as GPTScan Sun et al. (2024) and SMARTINV Wang et al. (2024), signify a monumental shift from traditional, pattern-based analysis methodologies towards more contextually aware and comprehensive inspection techniques. These cutting-edge tools extend their analytical prowess beyond static patterns by knitting together disparate threads of information, including the nuanced aspects of natural language documentation that detail the intended functions and transactional constructs of smart contracts.
@@
Expanding further on the key roles LLMs play, it’s worth noting the vast potential these models have in enhancing the entire lifecycle of smart contract development. From generation to maintenance, LLMs facilitate the crafting of more secure and robust smart contracts. They do so by potentially providing recommendations during the development phase, suggesting best practices, and even generating code snippets that align with security guidelines. Throughout the auditing process, tools like GPTScan and SMARTINV can continuously learn and adapt to new patterns of vulnerabilities emerging from the evolving landscape of blockchain technology and cyber threats. This dynamic learning process is pivotal, as it allows for the development of increasingly refined models capable of detecting even the most covert and sophisticated vulnerabilities.
@@ 3.2 LLM as Analyzers for abnormal transaction
The application of LLMs for blockchain transaction analysis Gai et al. (2023); Nicholls et al. (2023) underscores their crucial role in conducting real-time monitoring to detect signs of irregular or suspicious behavior. These models represent a significant advancement in the field, as they provide a more dynamic and adaptable approach to identifying potential threats within blockchain transactions.
@@ 3.3 LLM as Fuzzer for Smart Contract
Large Language Models (LLMs) have been increasingly employed to elevate the process of fuzzing, particularly in the realm of smart contract security analysis Shou et al. (2024); Zhang et al. (2024). This methodology involves utilizing LLMs to accurately assess the complexity and vulnerability likelihood of specific code regions within a smart contract. Consequently, these metrics serve to guide the direction and focus of fuzzers, steering them towards code segments that are more likely to harbor potential security threats.
The application of LLMs to fuzzing exercises significantly elevates the efficiency of these operations by narrowing down the vast state space that fuzzers typically navigate. This precision-targeted fuzzing approach contributes to higher coverage and reveals more vulnerabilities than conventional tools, especially those pertaining to the intricate nature of smart contract code that traditional methods may overlook.
@@ 3.4 LLM as Developer for Smart Contract
Recent studies Storhaug et al. (2023); Karanjai et al. (2023); Dade et al. (2023); Du and Tang (2024); Chen et al. (2023) have begun to scrutinize the efficacy and reliability of Large Language Models (LLMs) like ChatGPT and Google Palm2 in the automated generation of smart contracts. These smart contracts are integral to the blockchain ecosystem, executing agreements without the need for intermediaries, and their accuracy and security are paramount. The research primarily constructs a testing framework that assesses smart contracts on multiple fronts — validity, correctness, efficiency, security, and maintainability.
@@ 3.5 LLM as participants for Cryptocurrency community
Large Language Models (LLMs) such as GPT-3.5 and ChatGPT are emerging as powerful tools in the cryptocurrency community Trozze et al. (2023); Axelsen et al. (2023); Liu et al. (2024); Ziegler et al. (2024); wahidur2024enhancing, albeit with their respective strengths and weaknesses. Related works collectively depict a landscape where LLMs are being explored for their potential to revolutionize governance and legal processes within the high-stakes, highly volatile realm of cryptocurrency.
@@ 3.6 Miscellaneous
In addition, LLM is also used in blockchain security fields, involving smart contract compilers Karanjai et al. (2024), zero-knowledge proofs Wellington (2024), model training Luo et al. (2023), NFT generation He et al. (2023). We will introduce their applications in detail in future work.
@@ 4 Case study of LLM4BS
In this section, we engage in an in-depth examination through three distinct case studies, each serving to illustrate and shed light on the diverse and concrete applications of Large Language Models for Blockchain Systems (LLM4BS). These cases have been meticulously selected to encompass a broad range of scenarios.
@@ 4.1 LLM4Fuzz
LLM4FUZZ Shou et al. (2024) emerges as an innovative technique in the cybersecurity landscape, specifically in the niche of smart contract security within blockchain networks. It intricately combines the prowess of Large Language Models (LLMs) with fuzz testing methodologies to proactively unearth vulnerabilities that could potentially compromise the integrity of smart contracts.
@@ 4.2 SMARTINV
Proposed with the intention of enhancing the reliability and security of blockchain smart contracts, SMARTINV Wang et al. (2024) represents a significant breakthrough in the field. Its primary function is to infer invariants within smart contracts which can be integral in automating the process of identifying elusive bugs that typically elude conventional machine-auditing methods.
@@ 4.3 BLOCKGPT
BLOCKGPT Gai et al. (2023) serves as a paradigm shift in the domain of blockchain security, acting as a state-of-the-art Intrusion Detection System (IDS) specifically engineered to counteract and identify potentially malicious transactions within blockchain networks. The system is underpinned by a highly sophisticated large language model that has been meticulously trained with a significant corpus of transactional data from the Ethereum blockchain, one of the most widely utilized platforms in the industry.
@@ 5 Future Direction and Challenge of LLM4BS
Interdisciplinary Relationships: The essence of the next stage in LLM4BS is undeniably grounded in a harmonized interplay among the domains of artificial intelligence, cyber protection mechanisms, and distributed ledger technologies Kim et al. (2023); Meyer et al. (2023). This interdisciplinary collaboration is not merely additive but synergistic, as it draws upon the strengths and insights of each discipline to forge a formidable shield against cyber animosities. There is a clarion call within the academic and industrial spheres for a robust alliance, emphasizing that the amalgamation of cognitive computing with cryptographic resilience and decentralized architectures can lead to a paradigm shift in securing blockchain networks.
@@
Regulatory and Compliance Challenges: The shifting sands of regulatory frameworks demand not only compliance but a proactive engagement with regulatory bodies by scholars and practitioners in the LLM4BS field Wu et al. (2023); Teubner et al. (2023). This relationship is reciprocal; as regulatory agencies develop deeper understandings of the implications of integrating AI in blockchain, it is incumbent upon the actors within this space to advocate for regulations that encourage innovation while maintaining robust security measures. The dynamic interplay between cutting-edge technology and regulation is a delicate balance to strike, fostering a stable yet flexible platform for growth and adaptation in blockchain security solutions.
@@
Dynamic Security Threats: The cyber threat horizon is akin to a chimeric beast—constantly mutating and presenting unforeseen challenges Liu et al. (2023); Tan et al. (2023). Security models like LLM4BS must be engineered with inherent plasticity, allowing them to evolve alongside the threats they are designed to counteract. The integration of LLMs in blockchain security is not a static solution but a continually adapting safeguard, necessitating an expansive approach to cybersecurity that accounts for the proliferation of sophisticated cyberattacks as well as the subtleties of targeted breaches. Sustaining the integrity of blockchain transactions hinges on the preemptive identification and neutralization of these mercurial threats.
@@
Ethical Governance and Bias Mitigation: The ethical tapestry within which LLM4BS operates is rich and complex, mandating a conscientious approach towards the examination and resolution of security practices that may inadvertently propagate bias or unfair outcomes Aydin and Karaarslan (2023); Ray (2023). The quest for equitable algorithms expands beyond the technical realm, engaging with sociocultural dynamics and the moral dimensions of technological deployments. Therefore, a concerted effort in research that transcends statistical bias mitigation, touching upon philosophy, sociology, and ethics, is essential for fostering a climate where AI not only fortifies security but does so with an underlying commitment to justice and fairness.
@@
Energy Considerations and AI Sustainability: In addressing the carbon footprint of blockchain operations, there is also a pressing need to confront the energy-intensive nature of training and deploying Large Language Models Roumeliotis and Tselikas (2023); Qin et al. (2023); Miao et al. (2023). The ecological impact of these AI systems necessitates a dual strategy: enhancing algorithmic efficiency to reduce computational load and exploring alternative energy sources that can power these activities sustainably. This pursuit of ecological harmonization in the application of LLM4BS must be reflective of a broader commitment to sustainability across all aspects of blockchain technology, ensuring that the acceleration of security capabilities does not come at an unsustainable environmental cost.
@@ 6 Conclusion
In summing up this review on the incorporation of Large Language Models (LLMs) into blockchain security, we have traversed through the technological advancements and intricate challenges presented by this union. The potential of LLMs to augment security protocols in blockchain is clear, offering innovative solutions for smart contract auditing, identity verification, and anomaly detection. Yet, this potential comes with the necessity for vigilance regarding scalability, privacy, advancing cyber threats, and ethical implications of AI. As we look ahead, the success of LLMs in blockchain security depends not only on continuous technological refinement but also on ethical practices, regulatory alignment, and informed community engagement. The integration of LLM into blockchain heralds a transformative era in digital security that demands a collaborative approach, balancing innovation with prudent oversight to forge a resilient and equitable digital future.

### v3

@@ 3 Taxonomy of LLM4BS task
In this section, we introduce a thematic taxonomy devised to systematically categorize the body of literature about tasks associated with large language models for blockchain security (LLM4BS), emphasizing the function of the LLM within these contexts. Fig. 3 depicts the five applications of LLM4BS task, involving code audit of smart contracts §3.1, analysis of abnormal transactions §3.2, dynamic analysis of smart contracts §3.3, development of smart contracts §3.4, participants of cryptocurrency community §3.5, and other potential directions §3.6.
@@ 3.1 LLM as Code auditor on Smart Contracts
Table 2: Table of tools and models for code auditor
Research Functionality LLM(s) Publication
SMARTINV [24] Model finetuning approach Alpaca etc. SP,2024
GPTScan [13] Logic vulnerability detection GPT-3.5 etc. ICSE,2024
David et al. [25] LLM Audit Feasibility GPT-4 etc. arXiv,2023
Karanjai et al. [26] LLM code evaluation ChatGPT etc. BRAINS,2023
ContractArmor [27] Rule-based code analysis ChatGPT EUSPN,2024
Ortu et al. [28] Contract Automatic Repair ChatGPT etc. arXiv,2023
ASSBert [29] Contract vulnerability detection BERT JISA,2023
PSCVFinder [30] Contract vulnerability detection CodeT5 ISSRE,2023
LLM4Vuln [31] Vulnerability Reasoning Enhancement GPT-4 etc. arXiv,2024
TrustLLMf [32] Smart contracts audit CodeLlama-13b arXiv,2024
AuditGPT [33] ERC token audit ChatGPT etc. arXiv,2024
The application of LLM in the domain of smart contract code auditing and vulnerability detection can be succinctly encapsulated as follows: Advanced tools, such as SMARTINV [24], GPTScan [13], David et al. [25], Karanjai et al. [26], ContractArmor [27], Ortu et al. [28], ASSBert [29], PSCVFinder [30], LLM4Vuln [31], TrustLLMf [32] and AuditGPT [33]. The tools powered by Large Language Models in Table.2 signify a monumental shift from traditional, pattern-based analysis methodologies towards more contextually aware and comprehensive inspection techniques. These cutting-edge tools extend their analytical prowess beyond static patterns by knitting together disparate threads of information, including the nuanced aspects of natural language documentation that detail the intended functions and transactional constructs of smart contracts.
@@
Expanding further on the key roles LLMs play, it’s worth noting the vast potential these models have in enhancing the entire lifecycle of smart contract development [109]. From generation to maintenance, LLMs facilitate the crafting of more secure and robust smart contracts. They do so by potentially providing recommendations during the development phase, suggesting best practices, and even generating code snippets that align with security guidelines. Throughout the auditing process, tools like GPTScan and SMARTINV can continuously learn and adapt to new patterns of vulnerabilities emerging from the evolving landscape of blockchain technology and cyber threats. This dynamic learning process is pivotal, as it allows for the development of increasingly refined models capable of detecting even the most covert and sophisticated vulnerabilities.
@@ 3.2 LLM as Analyzers for abnormal transaction
Table 3: Table of tools and models for abnormal transaction analyzers
work function LLM(s) Publication
BLOCKGPT [21] Transaction anomaly detection Transformer arXiv,2023
Nicholls et al. [34] Logic vulnerability detection BERT etc. arXiv,2023
The application of LLMs for blockchain transaction analysis, such as BLOCKGPT [21] and Nicholls et al. [34], underscores their crucial role in conducting real-time monitoring to detect signs of irregular or suspicious behavior. These tools in Table 3 represents a significant advancement in the field, as they provide a more dynamic and adaptable approach to identifying potential threats within blockchain transactions.
@@ 3.3 LLM as Fuzzer for Smart Contract
Table 4: Table of tools and models for smart contract fuzzer
work function LLM(s) Publication
LLM4FUZZ [35] Fuzzing optimization tool Llama 2 arXiv,2024
ACFIX [36] AC vulnerability repair GPT-4 arXiv,2024
Large Language Models (LLMs) have been increasingly employed to elevate the process of fuzzing, particularly in the realm of smart contract security analysis ,such as LLM4FUZZ [35] and ACFIX [36]. This methodology in Table 4 involves utilizing LLMs to accurately assess the complexity and vulnerability likelihood of specific code regions within a smart contract. Consequently, these metrics serve to guide the direction and focus of fuzzers, steering them towards code segments that are more likely to harbor potential security threats.
The application of LLMs to fuzzing exercises significantly elevates the efficiency of these operations by narrowing down the vast state space that fuzzers typically navigate. This precision-targeted fuzzing approach contributes to higher coverage and reveals more vulnerabilities than conventional tools, especially those pertaining to the intricate nature of smart contract code that traditional methods may overlook [110].
@@ 3.4 LLM as Developer for Smart Contract
Table 5: Table of tools and models for smart contract developer
work function LLM(s) Publication
Storhaug et al. [22] Vulnerability-constrained decoding approach GPT-J-6B ISSRE,2023
karanjai et al. [26] Smart contract generation ChatGPT etc. BRAINS,2023
MazzumaGPT [37] Smart contract generation Davinci arXiv,2023
Du et al. [38] Audit capacity evaluation GPT-4 arXiv,2024
GPTutor [39] AI programming assistant GPT-3.5 etc. arXiv,2023
Recent studies in Table 5, such as Storhaug et al. [22], karanjai et al. [26], MazzumaGPT [37], Du et al. [38] and GPTutor [39], have begun to scrutinize the efficacy and reliability of Large Language Models (LLMs) like ChatGPT and Google Palm2 in the automated generation of smart contracts. These smart contracts are integral to the blockchain ecosystem, executing agreements without the need for intermediaries, and their accuracy and security are paramount. The research primarily constructs a testing framework that assesses smart contracts on multiple fronts — validity, correctness, efficiency, security, and maintainability.
@@ 3.5 LLM as participants for Cryptocurrency community
Table 6: Table of tools and models for Cryptocurrency community participants
work function LLM(s) Publication
Trozze et al. [40] Legal support tools GPT-3.5 etc. arXiv,2023
Axelsen et al. [41] Community moderation support ChatGPT arXiv,2023
Liu et al. [42] Blockchain-based Governance Framework   - IEEE Software,2024
ziegler et al. [43] Automating Contextual Classification GPT-4 arXiv,2024
GPTutor [44] Blockchain Revolutionizes Finance DistilBert etc. IEEE Access,2024
Large Language Models (LLMs) such as GPT-3.5 and ChatGPT are emerging as powerful tools in the cryptocurrency community, such as Trozze et al. [40], Axelsen et al. [41], Liu et al. [42], ziegler et al. [43] and GPTutor [44], albeit with their respective strengths and weaknesses. Related works in Table 6 collectively depict a landscape where LLMs are being explored for their potential to revolutionize governance and legal processes within the high-stakes, highly volatile realm of cryptocurrency.
@@ 3.6 Miscellaneous
Table 7: Table of other potential work
work function LLM(s) Publication
SolMover [45] Language Translation Framework Alpaca arXiv,2024
BasedAI [46] Privacy-Preserving Computation GPT-4 etc. arXiv,2024
BC4LLM [20] Secure Learning Path ChatGPT arXiv,2023
DLLM [47] Dynamic Language Modeling     - arXiv,2023
Diffusion-MVP [48] NFT creation platform Stable-Diffusion MM,2023
As we display in Table 7, LLM is also used in blockchain security fields, involving smart contract compilers [45], zero-knowledge proofs [46], model training [20, 47], NFT generation [48]. We will introduce their applications in detail in future work.
@@ 4 Case study of LLM4BS
Table 8: The table of the three cases on LLM4BS
Research Domain Publications Date faculty
LLM4FUZZ [35] Fuzz arxiv 2024 UC Berkeley
SMARTINV [24] program analysis IEEE S&P 2024 Columbia University
BLOCKGPT [21] transaction analysis arxiv 2023 University of California
In this section, we engage in an in-depth examination through three distinct case studies, each serving to illustrate and shed light on the diverse and concrete applications of Large Language Models for Blockchain Systems (LLM4BS). These cases in Table.8 have been meticulously selected to encompass a broad range of scenarios, comprising LLM4FUZZ [35] §4.1, SMARTINV [24] §4.2, BLOCKGPT [21] §4.3.
@@ 4.1 LLM4Fuzz
Figure 4: The architecture of LLM4FUZZ.
As we depicted in Fig.4, LLM4FUZZ [35] emerges as an innovative technique in the cybersecurity landscape, specifically in the niche of smart contract security within blockchain networks. It intricately combines the prowess of Large Language Models (LLMs) with fuzz testing methodologies to proactively unearth vulnerabilities that could potentially compromise the integrity of smart contracts.
@@ 4.2 SMARTINV
Proposed with the intention of enhancing the reliability and security of blockchain smart contracts, SMARTINV [24] represents a significant breakthrough in the field. Its primary function is to infer invariants within smart contracts which can be integral in automating the process of identifying elusive bugs that typically elude conventional machine-auditing methods. Fig.5 display the architecture of SMARTINV.
Figure 5: The architecture of SMARTINV.
@@ 4.3 BLOCKGPT
As we showed in Fig.6, BLOCKGPT [21] serves as a paradigm shift in the domain of blockchain security, acting as a state-of-the-art Intrusion Detection System (IDS) specifically engineered to counteract and identify potentially malicious transactions within blockchain networks. The system is underpinned by a highly sophisticated large language model that has been meticulously trained with a significant corpus of transactional data from the Ethereum blockchain, one of the most widely utilized platforms in the industry.
Figure 6: The architecture of blockgpt.
@@ 5 Future Direction and Challenge of LLM4BS
Interdisciplinary Relationships: The essence of the next stage in LLM4BS is undeniably grounded in a harmonized interplay among the domains of artificial intelligence, cyber protection mechanisms, and distributed ledger technologies [111, 112, 12]. This interdisciplinary collaboration is not merely additive but synergistic, as it draws upon the strengths and insights of each discipline to forge a formidable shield against cyber animosities. There is a clarion call within the academic and industrial spheres for a robust alliance, emphasizing that the amalgamation of cognitive computing with cryptographic resilience and decentralized architectures can lead to a paradigm shift in securing blockchain networks.
@@
Regulatory and Compliance Challenges: The shifting sands of regulatory frameworks demand not only compliance but a proactive engagement with regulatory bodies by scholars and practitioners in the LLM4BS field [5, 113]. This relationship is reciprocal; as regulatory agencies develop deeper understandings of the implications of integrating AI in blockchain, it is incumbent upon the actors within this space to advocate for regulations that encourage innovation while maintaining robust security measures. The dynamic interplay between cutting-edge technology and regulation is a delicate balance to strike, fostering a stable yet flexible platform for growth and adaptation in blockchain security solutions.
@@
Dynamic Security Threats: The cyber threat horizon is akin to a chimeric beast—constantly mutating and presenting unforeseen challenges [17, 114]. Security models like LLM4BS must be engineered with inherent plasticity, allowing them to evolve alongside the threats they are designed to counteract. The integration of LLMs in blockchain security is not a static solution but a continually adapting safeguard, necessitating an expansive approach to cybersecurity that accounts for the proliferation of sophisticated cyberattacks as well as the subtleties of targeted breaches. Sustaining the integrity of blockchain transactions hinges on the preemptive identification and neutralization of these mercurial threats.
@@
Ethical Governance and Bias Mitigation: The ethical tapestry within which LLM4BS operates is rich and complex, mandating a conscientious approach towards the examination and resolution of security practices that may inadvertently propagate bias or unfair outcomes [115, 116]. The quest for equitable algorithms expands beyond the technical realm, engaging with sociocultural dynamics and the moral dimensions of technological deployments. Therefore, a concerted effort in research that transcends statistical bias mitigation, touching upon philosophy, sociology, and ethics, is essential for fostering a climate where AI not only fortifies security but does so with an underlying commitment to justice and fairness.
@@
Energy Considerations and AI Sustainability: In addressing the carbon footprint of blockchain operations, there is also a pressing need to confront the energy-intensive nature of training and deploying Large Language Models [117, 118, 119]. The ecological impact of these AI systems necessitates a dual strategy: enhancing algorithmic efficiency to reduce computational load and exploring alternative energy sources that can power these activities sustainably. This pursuit of ecological harmonization in the application of LLM4BS must be reflective of a broader commitment to sustainability across all aspects of blockchain technology, ensuring that the acceleration of security capabilities does not come at an unsustainable environmental cost.
@@ 6 Conclusion
In summing up this review on the incorporation of Large Language Models (LLMs) into blockchain security, we have traversed through the technological advancements and intricate challenges presented by this combination of LLM4BS. The potential of LLMs to augment security protocols in the blockchain is clear, offering innovative solutions for smart contracts, detection of abnormal transactions, and development of the cryptocurrency community. Yet, this potential comes with the necessity for vigilance regarding scalability, privacy, advancing cyber threats, and ethical implications of AI. As we look ahead, the success of LLMs in blockchain security depends not only on continuous technological refinement but also on ethical practices, regulatory alignment, and informed community engagement. The integration of LLM into blockchain security heralds a transformative era that demands a collaborative approach, balancing innovation with prudent oversight to forge a resilient and equitable security future.